Theo is renting 2 types of tables for his party. One type seats 4 people and the other seats 6 people. If 36 will be at his party and he rents 7 tables, how many of each table does he rent?
Answer by amalm06(224)

The total number of people seated: 4x+6y=36
The total number of tables: x+y=7
Solve for x: x=7-y
4(7-y)+6y=36
28-4y+6y=36
2y=8
y=4
x=3
Therefore, he rents 3 tables seating 4 people and 4 tables seating 6 people (Answer)